# Information Limits of Joint Community Detection and Finite Group Synchronization

### v1

@@ 5.1 Preliminaries
    node $i\in S_{1}^{*}$ (resp. $S_{2}^{*}$) is connecting to more nodes in $S_{1}^{*}$ (resp. $S_{2}^{*}$) than in $S_{2}^{*}$ (resp. $S_{1}^{*}$) i.e.
    $|\mathcal{E}(i,S_{1}^{*}\backslash i)|>|\mathcal{E}(i,S_{2}^{*})|$
    (resp. $|\mathcal{E}(i,S_{2}^{*}\backslash i)|<|\mathcal{E}(i,S_{1}^{*})|$)
@@
    node $i\in H$ satisfies $|\mathcal{E}(i,S_{1}^{*}\backslash H)|>|\mathcal{E}(i,S_{2}^{*})|$
where $H$ is still the fixed subset of $S_{1}^{*}$ of size $\frac{n}{\gamma}$ and $\gamma=\log^{3}n$. In this way, Lemmas 5.4 and 5.5 still hold in this case. [^3] A slight difference is that we no longer need the event $F_{H}$, and therefore $E_{H}$ implies $E_{1}$ directly. Furthermore, we follow the steps in (55) and provide a lower bound on $\mathcal{P}\left(E_{H,A}^{(i)}\bigcap E_{B}^{(i,0)}\right)$ similar to Lemma 56. To this end, by using $r_{1}$ and $r_{2}$ defined in (53), we have
@@
$\log\mathcal{P}\left(E_{H,A}^{(i)}\bigcap E_{B}^{(i,0)}\right)$ $\geq\left(-\frac{a+b}{2}+\sqrt{\frac{ab}{M}}\right)\log n$ (59)
$\quad-o(\log n).$
@@
5.4 Proof of Theorem 3.3
To begin with, we follow the same path of proving Theorem 3.1 in Section 5.2 by writing out the probability that the MLE (27) fails to exactly recover the group elements as
@@
satisfies and therefore from Theorem 3.1, $\bm{\kappa}$ is incorrect w.p. at most $n^{-c}$. It remains to consider when $\bm{\kappa}=\bm{\kappa}^{*}$. In this case, the likelihood ratio in (31) satisfies
$\frac{\mathcal{P}(\bm{y}|\bm{x})}{\mathcal{P}(\bm{y}|\bm{x}^{*})}=\frac{\mathbbm{1}(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}^{*}_{\text{inner}})}{\mathbbm{1}(g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}^{*})}$
@@
$\mathcal{P}(E_{1})=\mathcal{P}(E_{2})\geq 1-n^{-c+o(1)}$
given the condition $\frac{a}{2}>1+c$ satisfies. Plugging (61) into (60) completes the proof.
5.5 Proof of Theorem 3.4
@@
$\frac{a+b}{2}-\sqrt{\frac{ab}{M}}<1$
satisfies, Theorem 3.2 suggests that the estimated community assignment $\bm{\kappa}$ is inaccurate with probability at least $3/5$. When this occurs, there exists at least one pair of wrong classified vertices $(i,j)$ such that $i\in S_{1}^{*}$ and $j\in S_{2}^{*}$ but $i\in S_{2}$ and $j\in S_{1}$. As a result, the estimated group elements on $i$ and $j$, denoted by $g_{i}$ and $g_{j}$, are completely determined by those edges across $S_{1}^{*}$ and $S_{2}^{*}$, whose group transformations are randomly drawn from $\mathcal{G}_{M}$. Therefore, the probability of having a wrong estimated $g_{i}$ or $g_{j}$ is at least $1-M^{-2}$. Combining this with the event of having a wrong community assignment $\bm{\kappa}$ yields the total probability of having the wrong group elements $\bm{g}$ at least $\frac{3}{5}\cdot(1-M^{-2})\geq 0.45$, where the equality holds when $M=2$. This completes the proof.
@@ Acknowledgments
Appendix A The giant component of Erdős–Rényi graphs
In this section, we briefly review several classical properties of Erdős–Rényi graph [20] that support our theory. Remarkably, most of the results in this section are by no means original, but we still provide our proofs in detail for completeness.
To begin with, given a network of size $n$, the Erdős–Rényi model denoted by $G(n,p)$  [^4] The model originally proposed by Erdős and Rényi is slightly different: instead of connecting each pair of vertices independently, it generates the graph by randomly assigning exactly $m$ edges to the network, and the resulting model is denoted by $G(n,m)$. However, the two models $G(n,m)$ and $G(n,p)$ are closely related, see [31, Chapter 1]. generates a random graph (denoted by $\mathscr{G}_{n}$) by connecting each pair of vertices independently with some probability $p$. In this section, our study focuses on the size of the largest connected component in $\mathscr{G}_{n}$, denoted by $Z_{n}$.
@@ Appendix A The giant component of Erdős–Rényi graphs
For completeness, a proof of Lemma A.1 is given in Appendix A.1, which is basically built from the analysis provided in [31]. Lemma A.1 indicates that the largest connected component will be the whole network w.h.p. when $a>1$. Furthermore, when $a<1$, it can be shown that the largest component will be giant w.h.p. such that
@@
To be concrete, [31, Theorem 2.14] states
$\mathcal{P}\left(Z_{n}>\left(1-n^{-a+\epsilon+o(1)}\right)n\right)>1-n^{-\epsilon}$ (62)
which holds for any $\epsilon<a$. (62) suggests that a giant component exists with a probability at least $n^{-\Theta(1)}$.
Unfortunately, (62) is too loose to be applied for our analysis in Section 5.2 and a tighter result is necessary. For example, we hope
$\mathcal{P}(Z_{n}>(1-o(1))n)>1-n^{-\omega(n)}$ (63)
@@
$\mathcal{P}\left(\beta_{1}\log n\leq Z_{n}\leq n-\frac{\beta_{2}n}{\log\log n}\right)\leq n^{-\Theta(\log n)}$ (64)
@@
Under the same setting of Lemma A.2, it satisfies
$\mathcal{P}\left(Z_{n}\leq\beta\log n\right)\leq n^{-\frac{an(1-o(1))}{2}}$ (65)
@@
Theorem A.4 (Giant component only when disconnected).
Under the same setting of Lemma A.2,
$\mathcal{P}\left(Z_{n}>n-\frac{\beta n}{\log\log n}\right)>1-n^{-\Theta(\log n)}$ (66)
@@
$>1-n^{-\Theta(\log n)}$
completes the proof. ∎
@@
$=\left(\frac{e^{k}}{k^{2}}\right)\cdot(a\log n)^{k-1}\cdot n^{-ak\left(1-\frac{k}{n}\right)+1}$ (69)
@@
$\mathbb{E}[r_{k}]\leq n^{-(1+c)k\left(1-\frac{k}{n}\right)(1-o(1))+1}$
@@
$\leq\mathcal{P}\left((r_{1}-\mathbb{E}[r_{1}])^{2}\geq\mathbb{E}^{2}[r_{1}]\right)\leq\frac{\text{Var}(r_{1})}{\mathbb{E}^{2}[r_{1}]}$ (71)
@@
$x_{i}=\begin{cases}1,&\quad\text{$i$ is isolated}\\ 0,&\quad\text{otherwise}\end{cases}$
@@
$\mathbb{E}[r_{1}]=\sum_{i=1}^{n}\mathbb{E}[x_{i}]=n(1-p)^{n-1}=n\cdot e^{-p(n-1)(1-o(1))}=n\cdot n^{a(1-o(1))}=n^{c-o(1)}$
@@
$\leq\sum_{\beta_{1}\log n\leq k\leq n-\frac{\beta_{2}n}{\log\log n}}\mathbb{E}[r_{k}]$ (72)
@@
$\leq-\Theta\left(\frac{n\log n}{\log\log n}\right)$
@@
$\mathcal{P}\left(\beta_{1}\log n\leq Z_{n}\leq n-\frac{\beta_{2}n}{\log\log n}\right)\leq n\cdot n^{-\Theta(\log n)}=n^{-\Theta(\log n)}$
@@
To interpret (75), first notice that
@@
assigns $kr_{k}$ nodes for those $r_{k}$ components of size $k$. Then, for each $kr_{k}$ nodes, $\binom{kr_{k}}{k,k,\ldots,k}$ further splits them into $r_{k}$ pieces. However, this overcounts the actual number by introducing the permutation over the $r_{k}$ pieces, therefore we divide it by $(r_{k})!$.
@@
$\leq\prod_{k=1}^{k_{\text{m}}}\left(k^{k-2}\cdot p^{k-1}\right)^{r_{k}}$
@@
=$\overset{(b)}{=}\frac{n(n-k_{\text{m}})}{2}$
@@
$\leq(1-p)^{m_{2,\bm{r}}}\cdot\prod_{k=1}^{k_{\text{m}}}\left(k^{k-2}\cdot p^{k-1}\right)^{r_{k}}$
@@
$\leq\sum_{\bm{r}}n!\cdot e^{-p\cdot m_{2,\bm{r}}}\cdot\left(\frac{(pk)^{n-\sum_{k=1}^{k_{\text{m}}}r_{k}}}{\prod_{k=1}^{k_{\text{m}}}(r_{k})!}\right)$ (76)
@@
$f(\bm{r})=\frac{n^{m}}{\prod_{k=1}^{k_{\text{m}}}(r_{k})!}\overset{(a)}{\leq}\frac{n^{m}}{\prod_{k=1}^{k_{\text{m}}}\frac{r_{k}^{r_{k}}}{e^{r_{k}-1}}}\leq n^{m}\Big{/}\underbrace{\prod_{k=1}^{k_{\text{m}}}\left(\frac{r_{k}}{e}\right)^{r_{k}}}_{=:g(\bm{r})}=\frac{n^{m}}{g(\bm{r})}$
@@
$r_{k}\geq 0,\;k=1,\ldots,k_{\text{m}}$
Notably, (80) ignores the constraint that $r_{k}\in\mathbb{Z}_{+}$ is an integer and instead extends to real number i.e. $r_{k}\in\mathbb{R}_{+}$, which significantly facilitates our analysis. The resulting optima is still a lower bound of $g(\bm{r})$ under $r_{k}\in\mathbb{Z}_{+}$.
@@
$r_{k}^{*}=c\cdot s^{k},\quad k=1,\ldots,k$ (82)
@@
$h(x):=\frac{1}{1-x}-\frac{k_{\text{m}}x^{k_{\text{m}}}}{1-x^{k_{\text{m}}}}$ (85)
@@
    $c^{\prime}(s)=\frac{(k_{\text{m}}+1)s^{k_{\text{m}}}-k_{\text{m}}s^{k_{\text{m}}+1}-1}{s^{2}(1-s^{k_{\text{m}}})^{2}}\leq 0$
@@ Appendix B Proof of lemmas
To begin with, recall the subgraph $G(S_{11})$ with size $|S_{11}|=\frac{n}{2}-\Delta=\Theta(n)$ follows Erdős–Rényi model. Then, let $z$ denote the size of the largest connected component of $G_{S_{11}}$, Theorem A.4 suggests a giant component in $G(S_{11})$ exists with high probability:
$\mathcal{P}\left(z\geq|S_{11}|-\frac{\beta n}{\log\log n}\right)\geq 1-n^{-\Theta(\log n)}$ (86)
@@
$E_{\mathrm{giant}}:\quad z\geq|S_{11}|-\frac{\beta n}{\log\log n}$
as the event of a giant component and $E_{\mathrm{giant}}^{c}$ as its complement. Then, $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1}\right)$ can be written as
@@
$\leq n^{-\Theta(\log n)}+\min\{1,\;\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1},E_{\mathrm{giant}}\right)\}$ (87)
@@
Now, we focus on bounding $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1},E_{\mathrm{giant}}\right)$ with a giant component exists. For each node $i\in S_{12}$, let us denote $r_{1,i}$ as the number of edges that it connects to $S_{11}$, and further, let $r_{1,\text{giant},i}\leq r_{1,i}$ denote the number of edges that connects to the giant component. Then, under a fixed set of $\left\{r_{1,i}\right\}_{i=1}^{\Delta}$ and $\left\{r_{1,\text{giant},i}\right\}_{i=1}^{\Delta}$, the event $E_{2,S_{1}}$ occurs with probability
@@
(88) can be obtained by using the same argument as in the proof of Lemma 5.1. That is, for each number of $r_{1,\text{giant},i}$ edges that connects to the giant component, it creates at least $r_{1,\text{giant},i}-1$ number of cycles, thus the corresponding probability that the cycle consistency satisfies is upper bounded by $M^{1-r_{1,\text{giant},i}}$. To proceed, we have
@@
$=\prod_{i=1}^{\Delta}\mathbb{E}\left[M^{1-r_{1,\text{giant},i}}\mid r_{1,i},E_{\text{giant}}\right]$ (89)
@@
Our next job is to upper bound $\mathbb{E}\left[M^{1-r_{1,\text{giant},i}}\mid r_{1,i},E_{\text{giant}}\right]$. First, recall that $z$ is denoted as the giant component size. We also define $z^{\prime}:=|S_{11}|-z$ as the size of the residual. Then, under a fixed $r_{1,i}$, the distribution of $r_{1,\text{giant},i}$ is given as
$\mathcal{P}\left(r_{1,\text{giant},i}=k\mid r_{1,i},E_{\text{giant}}\right)=\frac{\binom{z}{k}\cdot\binom{z^{\prime}}{r_{1,i}-k}}{\binom{|S_{11}|}{r_{1,i}}}$
@@
It is important to note that the summation over $k$ ranges from $\max\{0,\;r_{1,i}-z^{\prime}\}$ to $\min\{r_{1,i},\;z\}$ since there is at least $r_{1,i}-z^{\prime}$ edges connecting to the giant component when $r_{1,i}>z^{\prime}$, and also $r_{1,i}$ cannot exceed $z$. Then, in order to calculate (90), we consider two cases depending on the relation between $r_{1,i}$ and $z^{\prime}$:
@@
    $\leq\frac{Mz^{\prime}}{z\epsilon}(1+o(1))$ (92)
@@
    $\frac{f(k-1)}{f(k)}=O\left(\frac{\log\log\log n}{\log\log n}\right)=o(1)$
@@
$P_{\Delta}$ $\leq\min_{\alpha>0}\left(\sum_{r_{1}}\sum_{r_{2}}\sum_{r^{*}}\mathcal{P}_{\bm{x}^{*}}(r_{1})\mathcal{P}_{\bm{x}^{*}}(r_{2})\mathcal{P}_{\bm{x}^{*}}(r^{*})\cdot\left(\frac{Ma}{b}\right)^{\alpha(r_{1}+r_{2}-r^{*})}\cdot\left(n^{-\Theta(\log n)}+M^{2\Delta-(1-o(1))(r_{1}+r_{2})}\right)\right)$
$=\min_{\alpha>0}\left(\sum_{r}\sum_{r^{*}}\mathcal{P}_{\bm{x}^{*}}(r)\mathcal{P}_{\bm{x}^{*}}(r^{*})\cdot\left(\frac{Ma}{b}\right)^{\alpha(r-r^{*})}\cdot\left(n^{-\Theta(\log n)}+M^{2\Delta-(1-o(1))r}\right)\right)$
$=\min_{\alpha>0}\left(n^{-\Theta(\log n)}\cdot\mathbb{E}\left[\left(\frac{Ma}{b}\right)^{\alpha(r-r^{*})}\right]+M^{2\Delta}\cdot\mathbb{E}\left[\left(\frac{Ma}{b}\right)^{\alpha(r-r^{*})}\cdot M^{-(1-o(1))r}\right]\right).$ (94)

### v2

@@ 7.2 Case 2
    node $i\in S_{1}^{*}$ (resp. $S_{2}^{*}$) is connecting to more nodes in $S_{1}^{*}$ (resp. $S_{2}^{*}$) than in $S_{2}^{*}$ (resp. $S_{1}^{*}$) i.e., $|\mathcal{E}(i,S_{1}^{*}\backslash i)|>|\mathcal{E}(i,S_{2}^{*})|$ (resp. $|\mathcal{E}(i,S_{2}^{*}\backslash i)|<|\mathcal{E}(i,S_{1}^{*})|$),
@@
    node $i\in H$ satisfies $|\mathcal{E}(i,S_{1}^{*}\backslash H)|>|\mathcal{E}(i,S_{2}^{*})|$.
where $H$ is still the fixed subset of $S_{1}^{*}$ of size $\frac{n}{\gamma}$ and $\gamma=\log^{3}n$. In this way, Lemmas 7.1 and 7.2 still hold in this case. [^6] A slight difference is that we no longer need the event $F_{H}$, and therefore $E_{H}$ implies $E_{1}$ directly. Furthermore, we follow the steps in (64) and provide a lower bound on $\mathcal{P}\left(E_{H,A}^{(i)}\bigcap E_{B}^{(i,0)}\right)$ similar to Lemma 65. To this end, by using $r_{1}$ and $r_{2}$ defined in (62), we have
@@
8 Proof of Theorem 3.3
To begin with, we follow the same path of proving Theorem 3.1 in Section 6 by writing out the probability that the MLE (33) fails to exactly recover the group elements as
@@
is satisfied and therefore from Theorem 3.1, $\bm{\kappa}$ is incorrect w.p. at most $n^{-c}$. It remains to consider when $\bm{\kappa}=\bm{\kappa}^{*}$. In this case, the likelihood ratio in (37) satisfies
$\frac{\mathcal{P}(\bm{y}|\bm{x})}{\mathcal{P}(\bm{y}|\bm{x}^{*})}=\frac{\mathbbm{1}(g_{ij}=g_{i}g_{j}^{-1};\;\forall(i,j)\in\mathcal{E}^{*}_{\text{inner}})}{\mathbbm{1}(g_{ij}=g_{i}^{*}(g_{j}^{*})^{-1};\;\forall(i,j)\in\mathcal{E}_{\text{inner}}^{*})},$
@@
$\mathcal{P}(E_{1})=\mathcal{P}(E_{2})\geq 1-n^{-c+o(1)},$
given the condition $\frac{a}{2}>1+c$ is satisfied. Plugging (70) into (69) completes the proof.
9 Proof of Theorem 3.4
@@
$\frac{a+b}{2}-\sqrt{\frac{ab}{M}}<1,$
is satisfied, Theorem 3.2 suggests that the estimated community assignment $\bm{\kappa}$ is inaccurate with probability at least $3/5$. When this occurs, there exists at least one pair of wrong classified vertices $(i,j)$ such that $i\in S_{1}^{*}$ and $j\in S_{2}^{*}$ but $i\in S_{2}$ and $j\in S_{1}$. As a result, the estimated group elements on $i$ and $j$, denoted by $g_{i}$ and $g_{j}$, are completely determined by those edges across $S_{1}^{*}$ and $S_{2}^{*}$, whose group transformations are randomly drawn from $\mathcal{G}_{M}$. Therefore, the probability of having a wrong estimated $g_{i}$ or $g_{j}$ is at least $1-M^{-2}$. Combining this with the event of having a wrong community assignment $\bm{\kappa}$ yields the total probability of having the wrong group elements $\bm{g}$ at least $\frac{3}{5}\cdot(1-M^{-2})\geq 0.45$, where the equality holds when $M=2$. This completes the proof.
@@ Acknowledgments
Appendix A The giant connected component of Erdős–Rényi graphs
In this section, we briefly review several classical properties of the Erdős–Rényi graph [21] that support our theoretical framework. While most of the results presented are not original, we include detailed proofs for completeness. For readers interested in a comprehensive treatment of random graph theory, we refer to the excellent book by Frieze and Karoński [32].
To begin with, given a network of size $n$, the Erdős–Rényi model denoted by $G(n,p)$  [^7] The model originally proposed by Erdős and Rényi is slightly different: instead of connecting each pair of vertices independently, it generates the graph by randomly assigning exactly $m$ edges to the network, and the resulting model is denoted by $G(n,m)$. However, the two models $G(n,m)$ and $G(n,p)$ are closely related, see [32, Chapter 1]. generates a random graph (denoted by $\mathscr{G}_{n}$) by connecting each pair of vertices independently with some probability $p$. In this section, our study focuses on the size of the largest connected component in $\mathscr{G}_{n}$, denoted by $Z_{n}$.
@@ Appendix A The giant connected component of Erdős–Rényi graphs
For completeness, a proof of Lemma A.1 is given in Appendix A.1, which is basically built from the analysis provided in [32, Theorem 4.1]. Lemma A.1 indicates that the largest connected component will be the whole network w.h.p. when $a>1$. Furthermore, when $a<1$, it can be shown that the largest connected component will be giant w.h.p. such that
@@
To be concrete, by following standard techniques commonly used in the study of random graphs (e.g., [32, Theorem 2.14])888[32, Theorem 2.14] focuses on the regime where $p=\Theta(1/n)$, whereas our interest lies in the denser regime $p=\Theta(\log n/n)$. Therefore, we cannot directly apply [32, Theorem 2.14], but we adopt the same proof techniques to derive the bound in (26). In particular, we first show that there is no component of intermediate size (i.e., between $\beta_{1}\log n$ and $\beta_{2}n$ for some constants $\beta_{1},\beta_{2}>0$), and then we upper bound the total number of vertices in small components (those of size less than $\beta_{1}\log n$., we can derive the following result regarding $Z_{n}$,
$\mathcal{P}\left(Z_{n}>\left(1-n^{-a+\epsilon+o(1)}\right)n\right)>1-n^{-\epsilon},$ (71)
which holds for any $\epsilon<a$. (71) suggests that a giant connected component exists with a probability at least $n^{-\Theta(1)}$.
Unfortunately, (71) is too loose to be applied for our analysis in Section 6 and a tighter result is necessary. For example, we hope
$\mathcal{P}(Z_{n}>(1-o(1))n)>1-n^{-\omega(n)},$ (72)
@@
$\mathcal{P}\left(\beta_{1}\log n\leq Z_{n}\leq n-\frac{\beta_{2}n}{\log\log n}\right)\leq n^{-\Theta(\log n)},$ (73)
@@
$\mathcal{P}\left(Z_{n}\leq\beta\log n\right)\leq n^{-\frac{an(1-o(1))}{2}},$ (74)
@@
Theorem A.4 (Giant connected component).
$\mathcal{P}\left(Z_{n}>n-\frac{\beta n}{\log\log n}\right)>1-n^{-\Theta(\log n)},$ (75)
@@
$>1-n^{-\Theta(\log n)}.$
This completes the proof. ∎
@@
$=\left(\frac{e^{k}}{k^{2}}\right)\cdot(a\log n)^{k-1}\cdot n^{-ak\left(1-\frac{k}{n}\right)+1},$ (78)
@@
$\mathbb{E}[r_{k}]\leq n^{-(1+c)k\left(1-\frac{k}{n}\right)(1-o(1))+1},$
@@
$\leq\mathcal{P}\left((r_{1}-\mathbb{E}[r_{1}])^{2}\geq\mathbb{E}^{2}[r_{1}]\right)\leq\frac{\text{Var}(r_{1})}{\mathbb{E}^{2}[r_{1}]},$ (80)
@@
$x_{i}=\begin{cases}1,&\quad\text{$i$ is isolated}\\ 0,&\quad\text{otherwise},\end{cases}$
@@
$\mathbb{E}[r_{1}]=\sum_{i=1}^{n}\mathbb{E}[x_{i}]=n(1-p)^{n-1}=n\cdot e^{-p(n-1)(1-o(1))}=n\cdot n^{a(1-o(1))}=n^{c-o(1)},$
@@
$\leq\sum_{\beta_{1}\log n\leq k\leq n-\frac{\beta_{2}n}{\log\log n}}\mathbb{E}[r_{k}],$ (81)
@@
$\leq-\Theta\left(\frac{n\log n}{\log\log n}\right),$
@@
$\mathcal{P}\left(\beta_{1}\log n\leq Z_{n}\leq n-\frac{\beta_{2}n}{\log\log n}\right)\leq n\cdot n^{-\Theta(\log n)}=n^{-\Theta(\log n)},$
@@
To interpret (84), first notice that there are
@@
ways to assign $kr_{k}$ nodes for those $r_{k}$ components of size $k$. Then, for each $kr_{k}$ nodes, $\binom{kr_{k}}{k,k,\ldots,k}$ further splits them into $r_{k}$ pieces. However, this overcounts the actual number by introducing the permutation over the $r_{k}$ pieces, therefore we divide it by $(r_{k})!$.
@@
$\leq\prod_{k=1}^{k_{\text{m}}}\left(k^{k-2}\cdot p^{k-1}\right)^{r_{k}},$
@@
=$\overset{(b)}{=}\frac{n(n-k_{\text{m}})}{2},$
@@
$\leq(1-p)^{m_{2,\bm{r}}}\cdot\prod_{k=1}^{k_{\text{m}}}\left(k^{k-2}\cdot p^{k-1}\right)^{r_{k}},$
@@
$\leq\sum_{\bm{r}}n!\cdot e^{-p\cdot m_{2,\bm{r}}}\cdot\left(\frac{(pk)^{n-\sum_{k=1}^{k_{\text{m}}}r_{k}}}{\prod_{k=1}^{k_{\text{m}}}(r_{k})!}\right),$ (85)
@@
$f(\bm{r})=\frac{n^{m}}{\prod_{k=1}^{k_{\text{m}}}(r_{k})!}\overset{(a)}{\leq}\frac{n^{m}}{\prod_{k=1}^{k_{\text{m}}}\frac{r_{k}^{r_{k}}}{e^{r_{k}-1}}}\leq n^{m}\Big{/}\underbrace{\prod_{k=1}^{k_{\text{m}}}\left(\frac{r_{k}}{e}\right)^{r_{k}}}_{=:g(\bm{r})}=\frac{n^{m}}{g(\bm{r})},$
@@
$r_{k}\geq 0,\;k=1,\ldots,k_{\text{m}}.$
Notably, (89) ignores the constraint that $r_{k}\in\mathbb{Z}_{+}$ is an integer and instead extends to real number i.e.,$r_{k}\in\mathbb{R}_{+}$, which significantly facilitates our analysis. The resulting optima is still a lower bound of $g(\bm{r})$ under $r_{k}\in\mathbb{Z}_{+}$.
@@
$r_{k}^{*}=c\cdot s^{k},\quad k=1,\ldots,k_{\text{m}},$ (91)
@@
$h(x):=\frac{1}{1-x}-\frac{k_{\text{m}}x^{k_{\text{m}}}}{1-x^{k_{\text{m}}}},$ (94)
@@
    $c^{\prime}(s)=\frac{(k_{\text{m}}+1)s^{k_{\text{m}}}-k_{\text{m}}s^{k_{\text{m}}+1}-1}{s^{2}(1-s^{k_{\text{m}}})^{2}}\leq 0,$
@@ Appendix B Proof of lemmas
To begin with, recall the subgraph $G(S_{11})$ with size $|S_{11}|=\frac{n}{2}-\Delta=\Theta(n)$ follows Erdős–Rényi model. Then, let $z$ denote the size of the largest connected component of $G_{S_{11}}$, Theorem A.4 suggests a giant connected component in $G(S_{11})$ exists with high probability:
$\mathcal{P}\left(z\geq|S_{11}|-\frac{\beta n}{\log\log n}\right)\geq 1-n^{-\Theta(\log n)},$ (95)
@@
$E_{\mathrm{giant}}:\quad z\geq|S_{11}|-\frac{\beta n}{\log\log n},$
as the event of a giant connected component and $E_{\mathrm{giant}}^{c}$ as its complement. Then, $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1}\right)$ can be written as
@@
$\leq n^{-\Theta(\log n)}+\min\{1,\;\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1},E_{\mathrm{giant}}\right)\},$ (96)
@@
Now, we focus on bounding $\mathcal{P}_{\bm{x}^{*}}\left(E_{2,S_{1}}\mid r_{1},E_{\mathrm{giant}}\right)$ with a giant connected component exists. For each node $i\in S_{12}$, let us denote $r_{1,i}$ as the number of edges that it connects to $S_{11}$, and further, let $r_{1,\text{giant},i}\leq r_{1,i}$ denote the number of edges that are connected to the giant connected component. Then, under a fixed set of $\left\{r_{1,i}\right\}_{i=1}^{\Delta}$ and $\left\{r_{1,\text{giant},i}\right\}_{i=1}^{\Delta}$, the event $E_{2,S_{1}}$ occurs with probability
@@
(97) can be obtained by using the same argument as in the proof of Lemma 6.1. That is, for each number of $r_{1,\text{giant},i}$ edges that connects to the giant connected component, it creates at least $r_{1,\text{giant},i}-1$ number of cycles, thus the corresponding probability that the cycle consistency is satisfied is upper bounded by $M^{1-r_{1,\text{giant},i}}$. To proceed, we have
@@
$=\prod_{i=1}^{\Delta}\mathbb{E}\left[M^{1-r_{1,\text{giant},i}}\mid r_{1,i},E_{\text{giant}}\right],$ (98)
@@
Our next job is to upper bound $\mathbb{E}\left[M^{1-r_{1,\text{giant},i}}\mid r_{1,i},E_{\text{giant}}\right]$. First, recall that $z$ is denoted as the giant connected component size. We also define $z^{\prime}:=|S_{11}|-z$ as the size of the residual. Then, under a fixed $r_{1,i}$, the distribution of $r_{1,\text{giant},i}$ is given as
$\mathcal{P}\left(r_{1,\text{giant},i}=k\mid r_{1,i},E_{\text{giant}}\right)=\frac{\binom{z}{k}\cdot\binom{z^{\prime}}{r_{1,i}-k}}{\binom{|S_{11}|}{r_{1,i}}},$
@@
It is important to note that the summation over $k$ ranges from $\max\{0,\;r_{1,i}-z^{\prime}\}$ to $\min\{r_{1,i},\;z\}$ since there is at least $r_{1,i}-z^{\prime}$ edges connecting to the giant connected component when $r_{1,i}>z^{\prime}$, and also $r_{1,i}$ cannot exceed $z$. Then, in order to calculate (99), we consider two cases depending on the relation between $r_{1,i}$ and $z^{\prime}$:
@@
    $\leq\frac{Mz^{\prime}}{z\epsilon}(1+o(1)),$ (101)
@@
    $\frac{f(k-1)}{f(k)}=O\left(\frac{\log\log\log n}{\log\log n}\right)=o(1),$
@@
$P_{\Delta}$ $\leq\min_{\alpha>0}\left(\sum_{r_{1}}\sum_{r_{2}}\sum_{r^{*}}\mathcal{P}_{\bm{x}^{*}}(r_{1})\mathcal{P}_{\bm{x}^{*}}(r_{2})\mathcal{P}_{\bm{x}^{*}}(r^{*})\cdot\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha(r_{1}+r_{2}-r^{*})}\cdot\left(n^{-\Theta(\log n)}+M^{2\Delta-(1-o(1))(r_{1}+r_{2})}\right)\right)$
$=\min_{\alpha>0}\left(\sum_{r}\sum_{r^{*}}\mathcal{P}_{\bm{x}^{*}}(r)\mathcal{P}_{\bm{x}^{*}}(r^{*})\cdot\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha(r-r^{*})}\cdot\left(n^{-\Theta(\log n)}+M^{2\Delta-(1-o(1))r}\right)\right)$
$=\min_{\alpha>0}\left(n^{-\Theta(\log n)}\cdot\mathbb{E}\left[\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha(r-r^{*})}\right]+M^{2\Delta}\cdot\mathbb{E}\left[\left(\frac{Ma(1-q)}{b(1-p)}\right)^{\alpha(r-r^{*})}\cdot M^{-(1-o(1))r}\right]\right).$ (103)